Barnacle Industries was awarded a patent over 15 years ago for a unique industrial strength cleaner that removes barnacles and other particles from the hulls of ships. Thanks to its monopoly position, Barnacle has earned more than $160 million over the past decade. Its customers—spanning the gamut from cruise lines to freighters—use the product because it reduces their fuel bills. The annual (inverse) demand function for Barnacle’s product is given by P = 480 -0.00006Q, and Barnacle’s cost function is given by C(Q) = 390Q. Thanks to subsidies stemming from an energy bill passed by Congress nearly two decades ago, Barnacle does not have any fixed costs: The federal government essentially pays for the plant and capital equipment required to make this energy-saving product.

Absent this subsidy, Barnacle’s fixed costs would be about $9 million annually. Knowing that the company’s patent will soon expire, Marge, Barnacle’s manager, is concerned that entrants will qualify for the subsidy, enter the market, and produce a perfect substitute at an identical cost. With interest rates at 7 percent, Marge is considering a limit-pricing strategy.

What would Barnacle's profits be if Marge pursues a limit-pricing strategy if the subsidy is in place?

$  

What would be the profit of a new entrant if the subsidy is eliminated and Barnacle continues to produce the monopoly level of output?


$  
Answer #1

a) Total cost function of Bernacle when subsidy is in place,

TC= 390Q........(i)

Again interest rate = 7%

Therefore, new total cost of Bernacle with interest rate 7% is = 390Q*1.07= 417.3Q .......(ii)

Marginal cost function is arrived by taking derivative of total cost function (ii) with respect to Q

MC =417.3

Again the annual demand function of Bernacle is

P = 480-0.00006Q

and marginal revenue function is

MR = d(PQ)/dQ= 480-0.00012Q.......(ii)

Under limit price strategy, Bernacle will operate in perfect competitive market and under competitive equilibrium,

                  MR= MC

         or, 480-0.00012Q=417.3

       or, Q =522500

and P = 480-0.00006Q = 480-0.00006*522500= $448.65

Therefore, profits of Bernacle under competitive market equilibrium = Total revenue-Total cost= PQ-417.3Q

                                                            = 448.65*522500-417.3*522500= $16380375

b) If subsidy is eliminated, then total cost function of Bernacle is

TC= 9000000+390Q

Therefore, MC=390

For monopoly output level of Bernacle, MR=MC

In this case MR= 480-0.00012Q

   Therefore, 480-0.00012Q=390

   or, Q=750000 and P = 480-0.00006*750000= $435

The total cost of Bernacle when subsidy is eliminated = 9000000+390Q

Therefore, profits of Bernacle under monopoly condition = PQ-9000000-390Q

                                                               = 750000*435-9000000-390*435= $317080350